OpenStudy (secret-ninja):

What does this question even mean? I don't want the answer, I don't even know what it is saying. What does it want me to do? Q: Test the residuals of two other points to determine how well the line of best fit models the data.

OpenStudy (igreen):

I have no idea.

OpenStudy (secret-ninja):

Darn. Its the last question a long project I have for math, I've been on it for two days now, and all I need it to answer this. Can you tag any people that might know?

OpenStudy (igreen):

@hartnn @ganeshie8

ganeshie8 (ganeshie8):

whats your equation of best fit line ?

OpenStudy (secret-ninja):

y=1.07x+1

ganeshie8 (ganeshie8):

and you must be having some datapoint as well ? share complete details in the main question so that others get to know what exactly you're looking for

OpenStudy (secret-ninja):

Well, it is for a scatter plot, and I used the first point (58,59) and the last pint (73,74) to make the line of best fit. I don't know where to go from there.

ganeshie8 (ganeshie8):

Okay so what other points do you have ?

ganeshie8 (ganeshie8):

btw your best fit line is not matching with the points you have used

ganeshie8 (ganeshie8):

plugin (58, 59) and see if the equation satisfies

OpenStudy (secret-ninja):

I got 59=63

OpenStudy (secret-ninja):

I guess I am entirely wrong. :[

ganeshie8 (ganeshie8):

59 is never 63

OpenStudy (secret-ninja):

My equation should be y=1x+1 right?

OpenStudy (secret-ninja):

59=1x58+1

ganeshie8 (ganeshie8):

I dont know, the scatter plot is not precice enough to make out the points correctly

ganeshie8 (ganeshie8):

aren't you given them as ordered pairs ?

OpenStudy (secret-ninja):

Arm span is x and height is y

ganeshie8 (ganeshie8):

So you want to use below for best fit line, right ? `I used the first point (58,59) and the last pint (73,74)`

OpenStudy (secret-ninja):

Yep.

OpenStudy (secret-ninja):

aaah I wrote pint? I meant *point* >.<

OpenStudy (secret-ninja):

I think I got it sort of cleared up, but I still don't know what the question want me to do. :p

OpenStudy (igreen):

Oh, I did this assignment..I'm pretty sure they want you to use GeoGebra.

ganeshie8 (ganeshie8):

thats okay :) slope =1 is right

ganeshie8 (ganeshie8):

yeah `y = 1x + 1` looks good for best fit line!

OpenStudy (secret-ninja):

Now what do I do?

ganeshie8 (ganeshie8):

So you know that this line satisfies both first and last points accurately

ganeshie8 (ganeshie8):

so the first and last persons will be impressed with this best fit line as it correctly gives them their height if they just plugin the armspan

ganeshie8 (ganeshie8):

you need to see how accurately this best fit line matches the other people

ganeshie8 (ganeshie8):

pick any two ppoints from the table

OpenStudy (secret-ninja):

So i just get two random points that I have and prove it works?

OpenStudy (secret-ninja):

Oh okay. Gimme a sec to type.

ganeshie8 (ganeshie8):

it wont work for sure as it is just an "best approximation line"

ganeshie8 (ganeshie8):

you need to find out by how much error it fails

OpenStudy (igreen):

There's your line of best fit..

OpenStudy (secret-ninja):

70=1x69+1 70=69+1 70=70 68=1x67+1 68=67+1 68=68

ganeshie8 (ganeshie8):

hey pick some point that doesnt work :O

ganeshie8 (ganeshie8):

just for a variety

ganeshie8 (ganeshie8):

the points you picked are matching exactly to the line and so the error residue is 0 for them

ganeshie8 (ganeshie8):

that just means your best fit line works perfectly for those two points

ganeshie8 (ganeshie8):

try the point (62, 62)

OpenStudy (secret-ninja):

62=1x62+1 62=62+1 62=63

OpenStudy (secret-ninja):

Darn >.<

OpenStudy (secret-ninja):

I gtg. Bye.

OpenStudy (secret-ninja):

Thank for the help!

ganeshie8 (ganeshie8):

so the residue is 63-62 = 1 your best fit line gives an error of 1 inch or about 1.6% which is pretty good!